OpenStudy (anonymous):

If a triangle has an area 92 cm and a side length of 16 cm, what is the length of the altitude to this side of the triangle?

OpenStudy (raden):

Area = 1/2 * side * altitude 92 = 1/2 * 16 * altitude so, altitude = 92/(1/2 * 16) = .....